A point on the curve y = 6x - x2 at which the tangent to the curve is included at an angle of 45º to the line x + y = 0 is

Grade:

1 Answers

Arun
5 years ago
Dear student
 
dy/dx = 6- 2 x
 
now
 
|6 – 2 x| = -1 -1 / 2= -1
 
2 x = 7 hence x = 3.5 now you can calculate ordinate